A boy is 17 years old and his sister is twice as old. When the boy is 23 years old, what will be the age of his sister?
Found 2 solutions by Maths68, Hankmanscott:
Answer by Maths68(1474)

Age of boy = 17
Age of sister = 17*2 = 34
Boy will be 23 old in = 23-17 = 6 years
Age of sister will be in six year = 34+6 = 40 years.

Answer by Hankmanscott(3)